# Artificially intelligent agents in the social and behavioral sciences

### v1

@@ Abstract
We review the historical development and current trends of artificially intelligent agents (agentic AI) in the social and behavioral sciences—from the first programmable computers, and social simulations soon thereafter, to today’s experiments with large language models. This overview emphasizes the role of AI in the scientific process and the changes brought about, both through technological advancements and the broader evolution of science from around 1950 to the present. Some of the specific points we cover include: the challenges of presenting the first social simulation studies to a world unaware of computers, the rise of social systems science, intelligent game theoretic agents, the age of big data and the epistemic upheaval in its wake, and the current enthusiasm around applications of generative AI, and many other topics. A pervasive theme is how deeply entwined we are with the technologies we use to understand ourselves.
@@ 1 Introduction
Telling the story of AI and the social/behavioral sciences is challenging not only because of the entwined flows of ideas and influence, but also because of the many roles AI can have in it. First, as alluded to above, one could see AI as a building block of the scientific method, as an analysis tool, or as a digital twin technology (i.e., black box rendition of people or institutions), etc. Second, one could approach AI as an object of study and try to understand, for example, how behaviorism, which influenced Turing’s test [208, 24], transpires in LLMs and generative AI. Third, one could see the emergence of AI as a social phenomenon, where the imagination of investors, futurists, and scientists has aligned to bring us to where we are today. Finally, one could view AI as a social force changing both the interaction between people and the people themselves [224]. The availability of new research tools aside, to understand this emergent human-machine society is the most pressing motivation for experimenting with agentic AI at the time of writing [166, 158, 206, 207]. In Fig. 1, we sketch some approaches in the literature.
@@ 2 Simulations and early agent-based models
It might seem surprising at first, just how soon after the arrival of the first programmable computers, they were brought into use to simulate humans. An explanation is that the scientists behind these applications were, to a great extent, a part of the same intellectual movement that led to the development of computers themselves. By the time electronic computers became accessible for academic research, in the first few years of the 1950s, key ideas of simulational research were already public. For example, the idea that people’s roles can be characterized by how they are connected to others [140] and the thought that social agents could best be modeled as part random, part purpose-driven were established [167]. Furthermore, how to analyze the output from computer simulations was, of course, known from experimental and observational science. So even if the mode of scientific explanation manifested by computer simulations—famously described by James Coleman’s boat diagram [35, 225]—has taken a long time to become accepted by the mainstream, the elements were already there when the first computers became available to science. The transition was indeed so smooth that there were papers based on stochastic simulations with manual random number generation, like Helen Abbey’s study of the Reed-Frost epidemic model from 1952 [1]. [^1] For a more celebrated example, jumping ahead to 1969, stochastic modeling without the use of computers was also the premise for Thomas Schelling’s segregation model [180]. In the wake of the civil rights movement and the heightened racial tensions of the era, Schelling used a rule-based model of individuals moving their domiciles to make the point that people don’t have to have a strong racial bias for segregation to happen—it can emerge from the migration itself.
@@ 4.2 The early connectionists
5 Embodied cognition and autopoiesis
@@
Around the turn of the millennium, fields like robotics and human-computer interaction expanded into topics like the integration of socially aware agents in human populations (see, e.g., Ref. [38]). This stream of research was strongly influenced by embodied cognition. Furthermore, it approximated today’s discussion of AI in society, and even made some points that have since been lost. For example, this research stressed the importance of embodied perspectives and narrative understandings of courses of events if one seeks to make AI capable of replacing humans. This literature saw designing a robot friend beyond a simulacrum [23] as a distant goal, which, at the time of writing, is no longer on the horizon.
@@ 6.1 Agent-based models
Traditionally, the behavioral responses of agents have been encoded by simple decision rules [63, 62]. Although always considered unorthodox, ABMs have found their way into relatively conventional social science [117, 177]. A notable example is their use in analytical sociology [77, 78] and Epstein and Axtell’s more interdisciplinary Sugarscape simulations [44] (see Fig. 4). Economics has also seen a use of AMBs [46, 110], perhaps most frequently in the computational finance literature [108]. These studies include the use of reinforcement learning neural-network agents [18]. Finally, management science, which was keen to import complexity science concepts such as the “edge of chaos” [157], ventured into the realm of ABMs as well [147], especially to study innovation [111, 53] with the so-called NK model of evolution [99].
@@ 6.2 Game-theoretic agents
The use of ABMs in political science [9], economics [10], and, to some extent, anthropology [11] usually rests on a game-theoretical foundation [121] assuming people make decisions by maximizing utility functions. When different time horizons clash in the agents’ utility-optimization, we arrive at a social dilemma—a situation where the agents are faced with a choice to behave selfishly (typically leading to higher profits in a short perspective) or prosocially (engaging in cooperative efforts to build society). The study of human cooperation from this perspective has evolved in peculiar ways, where different sub-communities of the field have split and later merged again. In the 1980s, the inspiration came from population biology [123], with the sedentary transition of human prehistory as the explanandum [25]. In the early 1990s, the mathematics of complexity emerging from simplicity was a main driving force of the research [148]. In parallel to these developments, social dilemmas have been studied in economics in several waves. Notably, starting from Fehr et al.’s works on punishment and public goods [47], the early 2000s saw an increased interest in behavioral economics experiments about free riding and the over-exploitation of shared resources.
@@ 7 The age of big data
Around the first years of the new millennium, the AI winter gave way to a spring and summer that still lingers. The one-sentence summary of this development is: By using evermore resources, general methods can vastly outperform earlier, specialized ones [71]. The period up to today can be divided by the advent of near-human-level chatbots, where this section covers the first (and epistemically more disruptive) part. During this period, there was a concomitant advancement in hardware, data collection, and analysis methods, accompanied by growth in the information economy and an increased acceptance of the deployment of algorithmic decision making in areas such as insurance and corporate hiring [151]. This was fed into a feedback loop that catalyzed societal changes as interactions increasingly occurred online, thereby producing more data to fuel the algorithms. The zeitgeist is well captured by the McKinsey Global Institute’s 2011 report touting big data as “the next frontier for innovation, competition, and productivity” [119].
@@ 7.3 Computational social science and digital traces
The initially corporate term “data science” was soon adopted by academia, with universities launching degree programs, associations springing up, journals being founded, etc. Some authors see the field of computational social science as nacent of this development [217, 107, 175], whereas others point to deeper roots [91, 213]. Regardless, toward the turn of the millennium, an increasing availablility of data more or less directly recording human behavior—digital trace data—made the Internet a valuable data source for the social and behavior sciences [66].
@@
Figure 6: The DARPA Red Balloon challenge was a contest in social mobilization for geo-location to commemorate the Internet’s 40th anniversary. On October 29, 2009, ten red balloons were positioned at undisclosed locations across the continental USA (panel A). The first participant to correctly report the location of all was awarded \$40,000. The problem was designed to exploit (then) new technologies like social media and crowdsourcing to scale up the recruitment and organization of human work. Panel B illustrates the winning team’s strategy for incentivizing team-member recruitment inspired by Ref. [42]. Discovering a balloon earned a \$2000 share of the total prize sum, while recruiting someone earned half of that person’s award [199].
@@ 7.4 Crowdsourcing
Another data source (in addition to digital traces) that widespread online services made possible was crowdsourcing—the mobilization and coordination of people through digital platforms. Crowdsourcing has roots in optimization by competition, like Axelrod’s computer tournaments (to find the optimal strategy for the iterated Prisoner’s Dilemma [9]) and later data mining and machine learning contests like the KDD Cup and the Netflix Prize. It has been nicknamed “artificial artificial intelligence” for taking the distributed, computer-coordinated architecture from AI, but with humans rather than machines as computing nodes.
Another notable type of crowdsourcing challenge involved the planning and implementation of social mobilization. The most prominent example in this category was DARPA’s Red Balloon challenge, where teams attempted to locate ten balloons across the continental US. The winning team used a hierarchical system to distribute the award between people actually reporting the locations of the balloons and those recruiting the reporters [199, 161] (see Fig. 6).
An even more influential form of crowdsourcing-based science was the use of Amazon Mechanical Turk (MTurk) [26, 27]. Designed to recruit and pay workers for small computer-reportable tasks, originally motivated to build training data for image classification, it found use in the social and behavioral sciences as a platform for experiments. Even though MTurk experiments have always been criticized [100]—mostly because it is hard to know who the workers really are—they became very popular, with over a thousand papers per year by the mid-2010s [27].
@@ 7.5 Science of prediction
Fast forward to the present and this debate is still reverberating [86, 85], and, yes, there has been a revolution—especially in econometrics, but also in pockets of public health, political science, psychology, and sociology—but by the widespread adoption of causal inference methods [81], rather than validation by prediction. While this development surely has raised the awareness and quality of the scientific explanations coming out of these fields, some of the problems pointed out a decade earlier [215] remain but have fallen out of focus.
@@ 8.1 Technological developments
From 2017 onward, the AI research and industry witnessed an exponential growth driven predominantly by transformer-based models and generative AI. The Generative Pre-trained Transformer (GPT) series of models showed substantial improvements in natural language understanding and generation [164]. The Bidirectional Encoder Representations from Transformers (BERT) enabled contextual embeddings and thereby greatly advanced NLP tasks such as question answering and sentiment analysis [41]. These models are the basis for the LLMs that power the text generation in today’s AI agents.
@@ 8.2 Active topics
In an early paper of this kind, Park et al. demonstrated that communicating LLM agents could organize an event in collaboration and show other signs of emergent sociality [155] (see Fig. 7). Other papers replicated well-established human behavioral studies with AI: for example, Refs. [131, 50] studied a number of game theoretic experiments with LLMs to conclude that machines are slightly more cooperative than humans. Other studies, on the other hand, found them less cooperative [72]. This reflects the general sentiment at the time of writing: AI can emulate humans well enough to replace them in some intellectual activities, but is easily distinguishable from humans in certain tasks [207, 2, 200].
@@
If we are to believe the AI industry, we don’t have any alternative but a future with a much larger presence of AI. Understanding the emergent social consequences of this development is an important research direction [206, 158, 104, 184], and the motivation of several studies with AI agents.
Much of the research in this direction so far has focused on the ability of AI to influence people’s opinions. For example, Costello, Pennycook, and Rand [36] demonstrate that dialogues with AI can cause people to relinquish their beliefs in conspiracy theories for a considerable period. In another study, Hackenburg and Margetts [69] found that LLMs can effectively persuade people on political issues. With a careful study design, one could imagine studying AI-rich futures in the wild, but so far, we are only aware of ethically compromised attempts at this [150] (cf. Fig. 1D).
Another type of research on this theme concerns how human behavior changes when interacting or collaborating with machines compared to other humans. For example, Ref. [118] found that human-machine populations need both peer rewards and peer punishments to overcome human anti-machine bias in economic games.
@@
Foundation models like LLMs are capable of many forms of text and image analyses; maybe not as reliably or with the same interpretability as specialized NLP or image analysis tools, but very cheaply and accessibly.
For example, Bermejo et al. [19] note that for some classification tasks, LLMs can be more accurate than human annotators. Krugmann and Hartmann [103] are equally optimistic about LLMs’ capabilities for sentiment analysis [56], where Törnberg [202] gives a primer to social scientists for using LLM chatbots to analyze texts. Furthermore, Karjus [97] lists topic classification, historical event cause analysis, social network inference from literary texts, and missing data augmentation as some tasks AI is capable of aiding social scientists with.
@@ 8.3 The missing topic: Using AI to understand humans
Despite the surge of research with AI agents, and a flood of research-free perspective and opinion papers pointing out the possibilities [67, 195, 12], studies focusing on humans per se are bewilderingly rare. The majority of AI studies coming out today can, at best, be seen as pilot studies for future investigations of actual social behavior. Even agentic-AI papers whose stated motivation is to understand humans—like Argyle et al. [4], who use experiments with AI agents to validate theories about political persuasions—make the relative performance of various AI setups a take-home message. This is especially surprising given that interpreting human behavior, emotions, etc., to provide machines with an actionable input is a mature topic within the field of human-computer interaction [154]. Perhaps this fruit is just not as low-hanging as the other ones listed above and is simply biding its time to be picked.

### v2

@@ Abstract
We review the historical development and current trends of artificially intelligent agents (agentic AI) in the social and behavioral sciences—from the first programmable computers, and social simulations soon thereafter, to today’s experiments with large language models. This overview emphasizes the role of AI in the scientific process and the changes brought about, both through technological advancements and the broader evolution of science from around 1950 to the present. Some of the specific points we cover include: the challenges of presenting the first social simulation studies to a world unaware of computers, the rise of social systems science, intelligent game theoretic agents, the age of big data and the epistemic upheaval in its wake, and the current enthusiasm around applications of generative AI. A pervasive theme is how deeply entwined we are with the technologies we use to understand ourselves.
@@ 1 Introduction
Telling the story of AI and the social/behavioral sciences is challenging not only because of the entwined flows of ideas and influence, but also because of the many roles AI can have in it. First, as alluded to above, one could see AI as a building block of the scientific method, as an analysis tool, or as a digital twin technology (i.e., black box rendition of people or institutions). Second, one could approach AI as an object of study and try to understand, for example, how behaviorism, which influenced Turing’s test [224, 24], transpires in LLMs and generative AI. Third, one could see the emergence of AI as a social phenomenon, where the imagination of investors, futurists, and scientists has aligned to bring us to where we are today [152, 159]. Finally, one could view AI as a social force changing both the interaction between people and the people themselves [241]. The availability of new research tools aside, to understand this emergent human-machine society is the most pressing motivation for experimenting with agentic AI at the time of writing [181, 173, 222, 223]. In Fig. 1, we sketch some approaches in the literature.
@@ 2 Simulations and early agent-based models
When the first programmable computer, the ENIAC, began operating in the wake of World War II, it was naturally used for military purposes [76]. It is remarkable that only about five years later, the first social-science Ph.D. thesis based on computer science was presented [75]. This speed and ease of adoption of computer simulation can be explained in part by that many of the ideas behind electronic computers were well circulated at the time. For example, the thought that social agents could best be modeled as part random, part purpose-driven was already established [182]. How to analyze the output from computer simulations could, of course, be imported from experimental and observational science. The mode of scientific explanation manifested by computer simulations—expressed by, e.g., James Coleman’s boat diagram [38, 242]—was accepted when computers became available to scientific research. The transition was indeed so smooth that there were papers based on stochastic simulations with manual random number generation, like Helen Abbey’s study of the Reed-Frost epidemic model from 1952 [1].
@@ 4.2 The early connectionists
5 Neo cybernetics
In this section, we will cover some loosely connected streams of ideas that link the original cybernetics movement to the AI-infused, human-centric science of today. It contains more engineering-minded research, aiming to improve the world, compared to the following section.
5.1 Embodied cognition and autopoiesis
@@
5.2 Human-machine interaction
Although it is primarily concerned with advancing technology rather than our understanding of society and human behavior, the field of human-computer interaction (HCI) has much overlap with the themes in this paper. One influential opinion/framework paper in this area was the 1994 “Computers are social actors” [155], arguing that since people naturally anthropomorphize computers, it makes sense to approach HCI with methods from the behavioral and social sciences. Influenced by embodied cognition, this research expanded into topics like the integration of socially aware agents in human populations [41, 200]. Furthermore, it approximated today’s discussion of AI in society, and even made some points that have since been lost. For example, this research stressed the importance of embodied perspectives and narrative understandings of courses of events if one seeks to make AI capable of replacing humans.
@@ 6.1 Agent-based models
Traditionally, the behavioral responses of agents have been encoded by simple decision rules [67, 66]. Although always considered unorthodox, ABMs have found their way into relatively conventional social science [127, 192]. A notable example is their use in analytical sociology [83, 84] and Epstein and Axtell’s more interdisciplinary Sugarscape simulations [48] (see Fig. 4). Economics has also seen a use of ABMs [49, 120], perhaps most frequently in the computational finance literature [117]. These studies include the use of reinforcement learning neural-network agents [19]. Finally, management science, which was keen to import complexity science concepts such as the “edge of chaos” [172], ventured into the realm of ABMs as well [161], especially to study innovation [121, 57] with the so-called NK model of evolution [106].
@@ 6.2 Game-theoretic agents
The use of ABMs in political science [9], economics [10], and, to some extent, anthropology [11] usually rests on a game-theoretical foundation [132] assuming people make decisions by maximizing utility functions. When different time horizons clash in the agents’ utility-optimization, we arrive at a social dilemma—a situation where the agents are faced with a choice to behave selfishly (typically leading to higher profits in a short perspective) or prosocially (engaging in cooperative efforts to build society). The study of human cooperation from this perspective has evolved in peculiar ways, where different sub-communities of the field have split and later merged again. In the 1980s, the inspiration came from population biology [134], with the sedentary transition of human prehistory as the explanandum [25]. In the early 1990s, the mathematics of complexity emerging from simplicity was a main driving force of the research [162]. In parallel to these developments, social dilemmas have been studied in economics in several waves. Notably, starting from Fehr et al.’s works on punishment and public goods [51], the early 2000s saw an increased interest in behavioral economics experiments about free riding and the over-exploitation of shared resources. Some of these experiments have mixed human and bot populations, like Fig. 1C, usually with fairly simple rule-following bots [130]. These usually present ways bots can improve team performance [225]. For example, Shirado and Christakis show the ability of randomly behaving bots to push human participants out of local optima, toward better performance, in a color-coordination game [202].
@@ 7 The age of big data
Around the first years of the new millennium, the AI winter gave way to a spring and summer that still lingers. The one-sentence summary of this development is: By using evermore resources, general methods can vastly outperform earlier, specialized ones [77]. The period up to today can be divided by the advent of near-human-level chatbots, where this section covers the first (and epistemically more disruptive) part. During this period, there was a concomitant advancement in hardware, data collection, and analysis methods, accompanied by growth in the information economy and an increased acceptance of the deployment of algorithmic decision making in areas such as insurance, corporate hiring, and trading [166, 86]. This was fed into a feedback loop that catalyzed societal changes as interactions increasingly occurred online, thereby producing more data to fuel the algorithms. The zeitgeist is well captured by the McKinsey Global Institute’s 2011 report touting big data as “the next frontier for innovation, competition, and productivity” [129].
@@ 7.3 Computational social science and digital traces
The initially corporate term “data science” was soon adopted by academia, with universities launching degree programs, associations springing up, journals being founded, etc. Some authors see the field of computational social science as nacent of this development [234, 116, 190], whereas others point to deeper roots [98, 230]. Regardless, toward the turn of the millennium, an increasing availablility of data more or less directly recording human behavior—digital trace data—made the Internet a valuable data source for the social and behavior sciences [70], with social media as the chief data source [112, 35].
@@
Figure 6: The DARPA Red Balloon challenge was a contest in social mobilization for geo-location to commemorate the Internet’s 40th anniversary. On December 5, 2009, ten red balloons were positioned at undisclosed locations across the continental USA (panel A). The first participant to correctly report the location of all was awarded \$40,000. The problem was designed to exploit (then) new technologies like social media and crowdsourcing to scale up the recruitment and organization of human work. Panel B illustrates the winning team’s strategy for incentivizing team-member recruitment inspired by Ref. [46]. Discovering a balloon earned a \$2000 share of the total prize sum, while recruiting someone earned half of that person’s award [215].
@@ 7.4 Crowdsourcing
Another data source (in addition to digital traces) that widespread online services made possible was crowdsourcing—the mobilization and coordination of people through digital platforms. Crowdsourcing has roots in optimization by competition, like Axelrod’s computer tournaments (to find the optimal strategy for the iterated Prisoner’s Dilemma [9]) and later data mining and machine learning contests.
A notable type of crowdsourcing challenge involved the planning and implementation of social mobilization. The most prominent example in this category was DARPA’s Red Balloon challenge, where teams attempted to locate ten balloons across the continental US. The winning team used a hierarchical system to distribute the award between people actually reporting the locations of the balloons and those recruiting the reporters [215, 176] (see Fig. 6).
An even more influential form of crowdsourcing-based science was the use of Amazon Mechanical Turk (MTurk) [26, 27]. Designed to recruit and pay workers for small computer-reportable tasks, originally motivated to build training data for image classification, it found use in the social and behavioral sciences as a platform for experiments. Even though MTurk experiments have always been criticized [107]—mostly for uncontrolled biases due to the workforce’s anonymity—they became very popular, with over a thousand papers per year by the mid-2010s [27].
@@ 7.5 Science of prediction
Fast forward to the present and this debate is still reverberating [93, 92], and, yes, there has been a revolution—especially in econometrics, but also in pockets of public health, political science, psychology, and sociology—but by the widespread adoption of causal inference methods [88], rather than validating theories by their ability to predict. While this development surely has raised the awareness and quality of the scientific explanations coming out of these fields, some of the problems pointed out a decade earlier [232] remain but have fallen out of focus.
@@ 8.1 Technological developments
From 2017 onward, the AI research and industry witnessed a rapid growth driven predominantly by transformer-based models and generative AI. The Generative Pre-trained Transformer (GPT) series of models showed substantial improvements in natural language understanding and generation [179]. The Bidirectional Encoder Representations from Transformers (BERT) enabled contextual embeddings and thereby greatly advanced NLP tasks such as question answering and sentiment analysis [45]. These models are the basis for the LLMs that power the text generation in today’s AI agents.
@@ 8.2 Active topics
In an early paper of this kind, Park et al. demonstrated that communicating LLM agents could organize an event in collaboration and show other signs of emergent social behavior [170] (see Fig. 7). Other papers replicated well-established human behavioral studies with AI: for example, Refs. [142, 54] studied a number of game theoretic experiments with LLMs to conclude that machines are slightly more cooperative than humans. Other studies, on the other hand, found them less cooperative [78]. This reflects the general sentiment at the time of writing: AI can emulate humans well enough to replace them in some intellectual activities, but is easily distinguishable from humans in certain tasks [223, 2, 216].
@@
If we are to believe the AI industry, we don’t have any alternative but a future with a much larger presence of AI. Understanding the emergent social consequences of this development is an important research direction [222, 173, 113, 198, 31], and the motivation of several studies with AI agents.
A common topic on this topic has been to explore the ability of AI to influence people’s opinions. For example, Costello, Pennycook, and Rand [39] demonstrate that dialogues with AI can cause people to relinquish their beliefs in conspiracy theories for a considerable period. In another study, Hackenburg and Margetts [73] found that LLMs can effectively persuade people on political issues. With a careful study design, one could imagine studying AI-rich futures in the wild, but so far, we are only aware of ethically compromised attempts at this [165] (cf. Fig. 1D). As a final example, Tessler et al., found AI capable of
Another type of research on this theme concerns how human behavior changes when interacting or collaborating with machines compared to human-only groups. For example, Makovi et al. [128] found that human-machine populations need both peer rewards and peer punishments to overcome human anti-machine bias in economic games, whereas Köbis et al. [109] investigated dishonesty as a side-effect of delegating tasks to AI.
Finally, several studies have addressed the impact of AI on human creativity and productivity in intellectual team work. Some with positive conclusions, like Noy and Zhang [163] and Brynjolfsson et al. [30], both found workers more content and productive when performing writing tasks with AI assistance rather than without. Ref. [44] came to similar conclusions for AI-supported management consultancy. Conclusions from experiments with AI and creativity, on the other hand, seem task-dependent [119, 143].
@@
Foundation models like LLMs are capable of many forms of text and image analyses; maybe not as reliably or with the same interpretability as specialized NLP or image analysis tools, but very cheaply and accessibly. For example, Bermejo et al. [20] note that for some classification tasks, LLMs can be more accurate than human annotators. Krugmann and Hartmann [111] are equally optimistic about LLMs’ capabilities for sentiment analysis [60], while Refs. [218] and [17] presents the opportunities and risks of using LLMs to analyze text. Furthermore, Karjus [104] lists topic classification, historical event cause analysis, social network inference, and missing-data augmentation as suitable tasks for AI in the social sciences.
@@ 8.3 The missing topic: Using AI to understand humans
Despite the surge of research with AI agents, and a flood of research-free perspective and opinion papers pointing out the possibilities [71, 211, 12], studies focusing on humans per se are bewilderingly rare. The majority of AI studies coming out today can, at best, be seen as pilot studies for future investigations of actual social behavior. Even agentic-AI papers whose stated motivation is to understand humans—like Argyle et al. [4], who use experiments with AI agents to validate theories about political persuasions—make the relative performance of various AI setups a take-home message. This is especially surprising given that interpreting human behavior, emotions, etc., to provide machines with an actionable input is a mature topic within HCI [169]. Perhaps this fruit is just not as low-hanging as the other ones listed above and is simply biding its time to be picked.